OpenStudy (anonymous):

Find: integral (5e^x + 1/x)dx

OpenStudy (lgbasallote):

\(\LARGE \int \frac{5e^x + 1}{x} dx\)??

OpenStudy (anonymous):

\[\int\limits_{}^{}(5e ^{x}+1/x)dx\]

OpenStudy (unklerhaukus):

\[=\int 5e^x+\frac1 x \text d x=5e^x+\ln x+c\]

OpenStudy (anonymous):

@UnkleRhaukus are you sure of this answer?

OpenStudy (lgbasallote):

it's like \(\int 5e^x dx + \int \frac{1}{x} dx\)

OpenStudy (unklerhaukus):

the integral of e^x is e^x the integral of 1/x is ln|x|

OpenStudy (unklerhaukus):

you could put absolute value signs around the x in the logarithm \[=5e^x+\ln|x|+c \qquad c \in R\]